### v2

# Adversarial Observability and Performance Trade-offs in Optimal Control

Abstract
We develop a feedback controller that minimizes the observability of a set of adversarial sensors of a linear system, while adhering to strict closed-loop performance constraints. We quantify the effectiveness of adversarial sensors using the trace of their observability Gramian and its inverse, capturing both average observability and the least observable state directions of the system. We derive theoretical lower bounds on these metrics under performance constraints, characterizing the fundamental limits of observability reduction as a function of the performance trade-off. Finally, we show that the performance-constrained optimization of the Gramian’s trace can be formulated as a one-shot semidefinite program, while we address the optimization of its inverse through sequential semidefinite programming. Simulations on an aircraft show how the proposed scheme yields controllers that deteriorate adversarial observability while having near-optimal performance.
{IEEEkeywords}
Adversarial observability, optimal control, semidefinite programming.
1 Introduction
Autonomous systems often face adversaries who want to predict their current state or intent by deploying their own individual sensors. For example, in aviation, external observers can use aircraft position and velocity measurements obtained from radar to identify aircraft intent and predict future trajectories [25]. In space, external observers can determine a satellite’s active control mode using only unresolved optical images and without needing onboard telemetry data [8]. To ensure security, such systems must thus be able to maneuver in a way that minimizes the observability of potentially malicious observers.
A potential solution is moving target defense – a security method that aims to make the system’s evolution unpredictable and thus disrupts adversarial observability [16, 24, 12, 5]. The underlying concept is that, to impede an adversary from performing system reconnaissance or from predicting future trajectories, one should constantly change the system’s control mode. This induces a randomization in the evolution of the system that renders future tasks of the system difficult to identify. However, switching between different system modes can create discontinuities and transients in control execution, potentially degrading performance and compromising safety. Moreover, when the system has to commit to meeting a single objective, employing different control modes may not always be feasible. These observations highlight the need for observability-reduction mechanisms that do not rely on mode switching, but instead operate within a single, performance-constrained feedback architecture.
To obstruct the state reconstruction capability of adversarial sensors, several studies have investigated the design of a single controller that reduces observability [26, 27, 20, 1]. Related work has also studied the observability radius as a metric of distance to unobservability [4], network design for controllability metrics [2], and active defense strategies that mislead adversarial observers by modifying their available signals [22]. These approaches, however, either treat observability reduction as a binary constraint, focus on network topology rather than controller synthesis, or assume knowledge of a specific adversarial estimator. Moreover, they do not explicitly quantify the trade-off between closed-loop performance and observability reduction. This motivates feedback controllers that enable continuous, tunable trade-offs between performance and adversarial observability.
We consider a setup where the system’s operator, who has full state feedback, wants to design a linear controller to balance two conflicting objectives: i) to minimize the observability of a set of sensors which an adversary might be using to observe the system; and ii) to stabilize the system to the origin with optimal performance. We cast this dual-objective control problem as a constrained optimization of a metric of unobservability of the adversary’s sensors, subject to a trade-off hard constraint on the distance of closed-loop performance from optimality. In contrast to the aforementioned studies that treat unobservability as a binary property, we capture observability continuously, drawing from the sensor selection literature and leveraging observability Gramians [23, 21, 6, 10, 19].
Gramians are a prime tool for capturing the controllability and observability of linear systems in a continuous manner that goes beyond rank tests. Observability Gramians, in particular, must be inverted to perform least squares state estimation from partial observations [7], and hence their spectrum directly affects the quality of the resulting state estimate. For this reason, there is a rich body of literature that correlates how one should select the output nodes of a system with the underlying observability Gramian of those nodes [23, 21, 6, 10, 19]. Here, we explore this connection in the context of adversarial observability minimization. Unlike sensor selection, which chooses sensor configurations to maximize observability, our objective is to design the controller to actively reduce the adversary’s ability to observe the system.
We specifically quantify the observability of adversarial sensors by adopting the trace of the observability Gramian and its inverse. The former quantifies adversarial observability on average, across all observable directions, while the latter is skewed towards the less observable directions. For both of these metrics, we establish theoretical lower bounds on their optimal values as a function of the trade-off distance from optimal performance. We show that smaller performance trade-offs generally yield a sharper decrease in the trace of the observability Gramian, whereas larger performance trade-offs lead to a steeper increase in the trace of the Gramian’s inverse. Finally, we solve the constrained optimization of these metrics i) precisely for the trace of the Gramian, by proving it boils down to the solution of a semidefinite program (SDP); and ii) approximately for the trace of the Gramian inverse, through a sequential SDP algorithm that we obtain by applying the convex-concave procedure.
Notation: For any matrix $X$, $\left\lVert X\right\rVert$ denotes its operator norm. For a square matrix $X$, $\mathrm{tr}(X)$ denotes its trace, and $\underline{\lambda}(X)$, $\bar{\lambda}(X)$ denote its minimum and maximum eigenvalue when $X$ is symmetric, and $\kappa(X)$ the condition number. We denote $X\succ 0$ ($X\succeq 0$) if $X$ is positive definite (positive semidefinite), and $X\prec 0$ ($X\preceq 0)$ if $X$ is negative definite (negative semidefinite). We use $I$ to denote an identity matrix of appropriate dimensions.
2 Problem Formulation
Consider, for all $t\geq 0$, the continuous-time system
$\dot{x}(t)=Ax(t)+Bu(t),~x(0)=x_{0},~t\geq 0,$ (1)
where $x(t)\in\mathbb{R}^{n}$ is the state, $u(t)\in\mathbb{R}^{m}$ is the control input, and $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$ are the system’s state and input matrices.
2.1 Operator and Adversary Models
We assume that an adversary has access to the following partial observations of the state vector $x(t)$:
$y(t)=Cx(t).$
We will refer to this partial observation $y(t)\in\mathbb{R}^{p}$ as the output, and to $C\in\mathbb{R}^{p\times n}$ as the sensing matrix of the adversary. This sensing matrix corresponds to sensors that the adversary is personally employing to observe the state $x(t)$, but could also correspond to existing sensors of (1) that the adversary has compromised.
We also assume that the operator of system (1), who designs the control input $u(t)$, has access to the full state $x(t)$. In light of this sensing asymmetry, the operator is motivated to design a control law
$u(t)=Kx(t),$
$K\in\mathbb{R}^{m\times n}$, which does not only optimally stabilize system (1) to the origin, but also makes $x(t)$ difficult to observe from the output $y(t)$. In other words, the operator wants to design a feedback gain $K$ so that $A+BK$ is Hurwitz and nearly optimal, but also so that (A+BK,C)(A+BK,~C) is as unobservable as possible.
We consider the following assumptions, which describe the information structure of the operator and adversary.
Assumption 1.
$(A,B)$ is controllable.
Assumption 2.
The operator can measure the state $x(t)$ and knows the adversary’s matrix $C$. The adversary can measure the output $y(t)$ but cannot measure the operator’s control input $u(t)$.
Remark 1.
Assumption 1 is standard for the operator’s objective of optimally stabilizing (1) to the origin. Assumption 2 is indicative of the information asymmetry between the operator and the adversary, which the operator will use to make the task of observing $x(t)$ from $y(t)$ more difficult for the adversary. Knowledge of $C$ is realistic when the adversary’s sensing modality is constrained by physics (e.g., radar can only measure position and velocity [25], while optical telescopes capture external brightness but not onboard telemetry [8]). Finally, while the operator might often not know exactly which partial states the adversary can measure (as indicated by $C$), they are likely to know a superset of them, and hence Assumption 2 is not restrictive; we formalize this robustness in Proposition 2.
2.2 Operator’s Control Design Problem
Per the problem formulation, the operator considers two objectives when designing its control gain $K$: i) choosing $K$ to stabilize (1) to the origin optimally; and ii) choosing $K$ to make (A+BK,C)(A+BK,~C) as less observable as possible.
We capture the first of the two objectives with the cost function
$J_{s}(K)=\mathbb{E}\left[\int_{0}^{\infty}(x^{\mathrm{T}}(t)Qx(t)+u^{\mathrm{T}}(t)Ru(t))\mathrm{d}t\right],$ (2)
where the expectation is taken over the distribution of the initial condition $x_{0}\sim\mathcal{N}(0,V)$, and $V\succ 0$ is the initial condition’s covariance. Using standard linear systems theory, we can then simplify the expression of the cost (2) to
$J_{s}(K)=\mathrm{tr}(PV),$
where $P\succ 0$ is the unique solution of the Lyapunov equation
$(A+BK)^{\mathrm{T}}P+P(A+BK)+Q+K^{\mathrm{T}}RK=0.$ (3)
To capture the second objective of minimizing the observability of the pair (A+BK,C)(A+BK,~C), we resort to costs relating to observability Gramians. Following [21, 23], two relevant metrics are
$\begin{split}J_{o1}(K)&=\mathrm{tr}(WV),\\ J_{o2}(K)&=-\mathrm{tr}(W^{-1}V^{-1}),\end{split}$ (4)
where $W\succeq 0$ is the observability Gramian of the closed-loop pair (A+BK,C)(A+BK,~C), uniquely solving the Lyapunov equation
$(A+BK)^{\mathrm{T}}W+W(A+BK)+C^{\mathrm{T}}C=0.$ (5)
Remark 2.
The metrics (4) relate to the average energy the system releases through the sensors $C$. Specifically, $\mathbb{E}[\int_{0}^{\infty}y^{\mathrm{T}}(t)y(t)\textrm{d}t]=\mathbb{E}[\int_{0}^{\infty}x^{\mathrm{T}}(t)C^{\mathrm{T}}Cx(t)\textrm{d}t]=\mathrm{tr}(WV)=J_{o1}(K)$, whereas $J_{o2}(K)$ is motivated the inequality $\mathrm{tr}(W^{-1}V^{-1})\geq\frac{n^{2}}{\mathrm{tr}(WV)}$ and the fact that it becomes unbounded when the closed loop is not observable.
Remark 3.
Beyond the energy interpretation in Remark 2, the observability Gramian also governs estimation conditioning: in finite-horizon least-squares estimation, the sensitivity of the estimation error scales with the Gramian inverse [7], and more generally, directions that are weakly observable correspond to increased reconstruction uncertainty across estimation frameworks [6, 19]. An alternative measure of unobservability is the observability radius [4], which quantifies the smallest perturbation of $(A,C)$ that renders the system unobservable. However, our objective is not merely to certify proximity to rank loss, but to continuously degrade an adversary’s state-reconstruction capability under a hard performance constraint. Gramian-based metrics provide smooth measures of estimation conditioning [7], which makes them the appropriate choice in this setting.
Notably, to make the pair (A+BK,C)(A+BK,~C) as unobservable as possible, the operator would typically aim to either maximize $\mathrm{tr}(W^{-1}V^{-1})$ or minimize $\mathrm{tr}(WV)$. The difference between these two metrics lies in the aspect of observability they emphasize. The quantity $\mathrm{tr}(W^{-1}V^{-1})$ is better suited for capturing how close the pair (A+BK,C)(A+BK,~C) is to being unobservable, as it becomes unbounded (infinite) if the system is not observable. This makes it particularly effective when the objective is to suppress the most weakly observable modes. On the other hand, $\mathrm{tr}(WV)$ is better suited for quantifying observability on average, across all directions of the state space. It remains finite even when (A+BK,C)(A+BK,~C) is unobservable, and is thus a preferable metric when one does not place much emphasis on complete unobservability, or when a well-defined metric is needed in both observable and unobservable regimes.
In view of the discussion above, to co-optimize the performance of the closed-loop plant matrix $A+BK$ as well as the unobservability of the pair (A+BK,C)(A+BK,~C), we focus on two optimization problems. The first problem minimizes the first unobservability objective in (4), subject to a constraint that limits the degradation in performance, as quantified by the cost (2) and a trade-off parameter $\lambda>0$. Specifically, this constraint is $J_{s}(K)-J_{s}(K^{\star})\leq\lambda$, or equivalently
$\mathrm{tr}((P-P^{\star})V)\leq\lambda,$
where $K^{\star}=-R^{-1}B^{\mathrm{T}}P^{\star}$, and $P^{\star}\succ 0$ is the positive definite solution of the algebraic Riccati equation
$A^{\mathrm{T}}P^{\star}+P^{\star}A+Q-P^{\star}BR^{-1}B^{\mathrm{T}}P^{\star}=0.$ (6)
We summarize the overall problem as follows.
Problem 1.
Solve the optimization problem:
$\begin{split}\min_{K,P,W}\quad&\mathrm{tr}(WV)\\ \textrm{s.t.}\quad&(A+BK)^{\mathrm{T}}P+P(A+BK)+Q+K^{\mathrm{T}}RK=0,\\ &(A+BK)^{\mathrm{T}}W+W(A+BK)+C^{\mathrm{T}}C=0,\\ &\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~W\succeq 0.\end{split}$
The second problem of co-optimizing (2) and the second observability metric in (4) is more intricate. The main reason is that when the pair $(A+BK,C)$ is not observable, the quantity $\mathrm{tr}(W^{-1}V^{-1})$ becomes undefined due to the observability Gramian $W$ being singular. This technicality commonly appears in controllability and observability quantification problems, and is bypassed by regularizing the observability Gramian with a small positive definite matrix [13]. Hereon, we adopt this standard regularization approach, but instead of applying the regularization directly to $W$ after its computation, we apply it at the level of the Lyapunov equation (5), from which $W$ is obtained. As we will see later, this choice is crucial for obtaining convergent algorithmic solutions. We thus obtain the following alternative optimization for designing $K$.
Problem 2.
Select $\epsilon>0$ and solve the optimization problem:
$\begin{split}\min_{K,P,W_{\epsilon}}\quad&-\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})\\ \textrm{s.t.}\quad&(A+BK)^{\mathrm{T}}P+P(A+BK)+Q+K^{\mathrm{T}}RK=0,\\ &(A+BK)^{\mathrm{T}}W_{\epsilon}+W_{\epsilon}(A+BK)+C^{\mathrm{T}}C+\epsilon I=0,\\ &\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~W_{\epsilon}\succeq 0.\end{split}$
Remark 4.
Per [13], one should select $\epsilon$ to be close to 0.
Both Problem 1 and 2 involve bilinear constraints. Problem 2 is further complicated by its nonlinear cost function. We will provide lower bounds to their values that showcase the fundamental limits of the trade-off between adversarial observability and performance, as it relates to $\lambda$. Moreover, we will derive an algorithm that exactly solves Problem 1 using an equivalent SDP, and an algorithm that approximately solves Problem 2 through sequential SDP.
3 Adversarial Observability and Performance Trade-offs
In this section, we study the adversarial observability and performance trade-off in Problems 1 and 2. We specifically provide lower bounds to the values of these problems as a function of the trade-off parameter $\lambda$, and study their robustness to uncertainty in the adversary’s sensing matrix.
3.1 Quantitative Lower Bounds
Define the matrices $Z^{\star},~S^{\star},~U^{\star}\succ 0$ that solve
$(A+BK^{\star})Z^{\star}+Z^{\star}(A+BK^{\star})^{\mathrm{T}}+V=0,$ (7)
$(A+BK^{\star})S^{\star}+S^{\star}(A+BK^{\star})^{\mathrm{T}}+I=0,$ (8)
$(A+BK^{\star})^{\mathrm{T}}U^{\star}+U^{\star}(A+BK^{\star})+I=0.$ (9)
We present an intermediate result that bounds the norm of the gain distance $\tilde{K}:=K-K^{\star}$, where $K$ is the solution to either Problem 1 or 2, and $K^{\star}$ is the optimal gain.
Lemma 1.
It holds that $\|\tilde{K}\|\leq f(\lambda)$, where
$f(\lambda):=\frac{\lambda\|Z^{\star}\|\|{V^{-1}}\|\left\lVert B\right\rVert}{\underline{\lambda}(Z^{\star})\underline{\lambda}(R)}\left(1+\sqrt{1+\frac{\underline{\lambda}(Z^{\star})\underline{\lambda}(R)}{\lambda\|Z^{\star}\|^{2}\|{V^{-1}}\|^{2}\left\lVert B\right\rVert^{2}}}\right).$
Proof.
From first constraint of Problem 1 or 2 and from (6):
$(A+BK)^{\mathrm{T}}P+P(A+BK)+Q+K^{\mathrm{T}}RK=0,$
$(A+BK)^{\mathrm{T}}P^{\star}+P^{\star}(A+BK)$
$\qquad\qquad\qquad+Q+K^{\star^{\mathrm{T}}}RK^{\star}-\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}P^{\star}-P^{\star}B\tilde{K}=0.$
Subtracting these two equations, using the property $B^{\mathrm{T}}P^{\star}=-RK^{\star}$, and defining $\tilde{P}=P-P^{\star}$, we get
$(A+BK)^{\mathrm{T}}\tilde{P}+\tilde{P}(A+BK)+\tilde{K}^{\mathrm{T}}R\tilde{K}=0,$
and, by adding and subtracting identical terms:
$(A+BK^{\star})^{\mathrm{T}}\tilde{P}+\tilde{P}(A+BK^{\star})+\tilde{K}^{\mathrm{T}}R\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K}=0.$
The implicit solution to this equation is given by:
$\tilde{P}=\int_{0}^{\infty}e^{(A+BK^{\star})^{\mathrm{T}}t}(\tilde{K}^{\mathrm{T}}R\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K})e^{(A+BK^{\star})t}\textrm{d}t.$
Using the constraint $\mathrm{tr}((P-P^{\star})V)=\mathrm{tr}(\tilde{P}V)\leq\lambda$, this implies
$\mathrm{tr}(\tilde{P}V)$ $=\mathrm{tr}\Big({\int_{0}^{\infty}}e^{(A+BK^{\star})^{\mathrm{T}}t}(\tilde{K}^{\mathrm{T}}R\tilde{K}$
$\qquad\qquad\qquad+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K})e^{(A+BK^{\star})t}\textrm{d}tV\Big)$
$=\mathrm{tr}\Big({\int_{0}^{\infty}}e^{(A+BK^{\star})t}Ve^{(A+BK^{\star})^{\mathrm{T}}t}\textrm{d}t(\tilde{K}^{\mathrm{T}}R\tilde{K}$
$)\displaystyle\qquad\qquad\qquad+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K}$$\Big)$
$)\displaystyle=\mathrm{tr}(Z^{\star}(\tilde{K}^{\mathrm{T}}R\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K}$))
$=\mathrm{tr}(Z^{\star}\tilde{K}^{\mathrm{T}}R\tilde{K})+2\mathrm{tr}(\tilde{P}B\tilde{K}Z^{\star})\leq\lambda,$ (10)
where $Z^{\star}$ solves the Lyapunov equation (7). Note that since $P^{\star}$ is the ARE solution (6), it follows that $P^{\star}\preceq P$, hence $\tilde{P}\succeq 0$. Therefore, using Fact 8.12.28 from [3], (10) yields
$\mathrm{tr}(Z^{\star}\tilde{K}^{\mathrm{T}}R\tilde{K})-2\mathrm{tr}(\tilde{P})\|B\tilde{K}Z^{\star}\|\leq\lambda.$ (11)
Moreover, we have
$\mathrm{tr}(\tilde{P})=\mathrm{tr}(V^{1/2}\tilde{P}V^{1/2}V^{-1})\leq\mathrm{tr}(\tilde{P}V)\|V^{-1}\|\leq\lambda\|V^{-1}\|.$
Hence, (11) yields
$\underline{\lambda}(Z^{\star})\underline{\lambda}(R)\|\tilde{K}\|^{2}-2\lambda\|Z^{\star}\|\|{V^{-1}}\|\left\lVert B\right\rVert\|\tilde{K}\|-\lambda\leq 0.$
Solving this quadratic inequality we obtain
$\|\tilde{K}\|$ $\leq\frac{\lambda\|Z^{\star}\|\|{V^{-1}}\|\left\lVert B\right\rVert}{\underline{\lambda}(Z^{\star})\underline{\lambda}(R)}\left(1+\sqrt{1+\frac{\underline{\lambda}(Z^{\star})\underline{\lambda}(R)}{\lambda\|Z^{\star}\|^{2}\|{V^{-1}}\|^{2}\left\lVert B\right\rVert^{2}}}\right),$
which is the required result.
Next, denote the value of Problem 1 with respect to $\lambda$ as $J_{1}(\lambda)$, and that of Problem 2 as $J_{2}(\lambda)$. Using Lemma 1, we provide lower bounds to the values of these problems as functions of the observability-performance trade-off parameter $\lambda$.
Theorem 1.
The following hold true.
  1. 1.
    For all $\lambda\geq 0$:
    $J_{1}(\lambda)\geq\frac{1}{1+2\mathrm{tr}(Z^{\star})f(\lambda)\|V^{-1}\|\left\lVert B\right\rVert}J_{1}(0),$ (12)
  2. 2.
    There exists $\lambda^{\star}>0$ such that for all $\lambda\leq\lambda^{\star}$
    $J_{2}(\lambda)\geq\frac{1}{1-\frac{2\kappa(V)f(\lambda)\left\lVert V\right\rVert\left\lVert B\right\rVert\|U^{\star}\|\mathrm{tr}(W_{\epsilon}^{0})\mathrm{tr}((W^{0}_{\epsilon})^{-1})}{1-2\mathrm{tr}(S^{\star})f(\lambda)\|B\|}}J_{2}(0),$ (13)
    where $W_{\epsilon}^{0}$ corresponds to $W_{\epsilon}$ under $K=K^{\star}$.
Proof.
To prove item 1, the second constraint of Problem 1 yields
$(A+BK^{\star})^{\mathrm{T}}W+W(A+BK^{\star})+C^{\mathrm{T}}C+WB\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}W=0.$
Therefore, denoting the Gramian in Problem 1 resulting with $\lambda=0$ as $W^{0}$, we obtain
$W=W^{0}+\int_{0}^{\infty}e^{(A+BK^{\star})^{\mathrm{T}}t}(WB\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}W)e^{(A+BK^{\star})t}\mathrm{d}t.$
Applying traces to each side of this equality:
$\mathrm{tr}$ $(WV)=\mathrm{tr}(W^{0}V)$
$+\mathrm{tr}\left(\int_{0}^{\infty}e^{(A+BK^{\star})t}Ve^{(A+BK^{\star})^{\mathrm{T}}t}\textrm{d}t(WB\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}W)\right)$
$=\mathrm{tr}(W^{0}V)+\mathrm{tr}(Z^{\star}(WB\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}W)).$ (14)
Using Fact 8.12.28 from [3] and Lemma 1
$\mathrm{tr}(WV)$ $\geq\mathrm{tr}(W^{0}V)-2\mathrm{tr}(Z^{\star})\left\lVert W\right\rVert\left\lVert B\right\rVert\|\tilde{K}\|$
$\geq\mathrm{tr}(W^{0}V)-2\mathrm{tr}(Z^{\star})f(\lambda)\mathrm{tr}(WV)\|V^{-1}\|\left\lVert B\right\rVert.$
Rearranging terms, we get $\mathrm{tr}(WV)\geq\mathrm{tr}(W^{0}V)/(1+2\mathrm{tr}(Z^{\star})f(\lambda)\|V^{-1}\|\left\lVert B\right\rVert)$, which proves item 1.
To prove item 2, because of the second constraint of Problem 2, we have similarly that $W_{\epsilon}=W^{0}_{\epsilon}+X,$ where $W_{\epsilon}^{0}$ denotes the Gramian in Problem 2 for $\lambda=0$, and $X=\int_{0}^{\infty}e^{(A+BK^{\star})^{\mathrm{T}}t}(W_{\epsilon}B\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}W_{\epsilon})e^{(A+BK^{\star})t}\mathrm{d}t$. Using the Woodbury identity [14] we have $W_{\epsilon}^{-1}V^{-1}=(W^{0}_{\epsilon})^{-1}V^{-1}-(W^{0}_{\epsilon})^{-1}XW_{\epsilon}^{-1}V^{-1}$, hence
$\mathrm{tr}(W_{\epsilon}^{-1}$ $V^{-1})=\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1})-\mathrm{tr}((W^{0}_{\epsilon})^{-1}XW_{\epsilon}^{-1}V^{-1})$
$\leq\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1})+\kappa(V)\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1})\|XW_{\epsilon}^{-1}\|$
$\leq(1+\kappa(V)\|W_{\epsilon}^{-1}\|\left\lVert X\right\rVert)\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1}).$ (15)
Moreover, by the definition of $X$, we have $X\preceq 2\|\tilde{K}\|\left\lVert B\right\rVert\left\lVert W_{\epsilon}\right\rVert U^{\star}\preceq 2f(\lambda)\left\lVert B\right\rVert\left\lVert W_{\epsilon}\right\rVert U^{\star}$. Combining this with (15) yields
$\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})$
$\leq(1{+}2\kappa(V)f(\lambda)\left\lVert B\right\rVert\left\lVert W_{\epsilon}\right\rVert\|W_{\epsilon}^{-1}\|\|U^{\star}\|)\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1})$
$\leq(1+2\kappa(V)f(\lambda)\left\lVert B\right\rVert\mathrm{tr}(W_{\epsilon})\mathrm{tr}(W_{\epsilon}^{-1})\|U^{\star}\|)\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1}).$
Regrouping terms:
$\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})\\ \leq\frac{\mathrm{tr}((W_{\epsilon}^{0})^{-1}V^{-1})}{1-2\kappa(V)f(\lambda)\left\lVert V\right\rVert\left\lVert B\right\rVert\|U^{\star}\|\mathrm{tr}(W_{\epsilon})\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1})}.$ (16)
Finally, similar to (14) in item 1 we have
$\mathrm{tr}(W_{\epsilon})$ $=\mathrm{tr}(W^{0}_{\epsilon})+\mathrm{tr}(S^{\star}(W_{\epsilon}B\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}W_{\epsilon})))$
$\leq\mathrm{tr}(W^{0}_{\epsilon})+2\mathrm{tr}(S^{\star})\|W_{\epsilon}B\tilde{K}\|$
$\leq\mathrm{tr}(W^{0}_{\epsilon})+2\mathrm{tr}(S^{\star})f(\lambda)\|B\|\mathrm{tr}(W_{\epsilon}),$
where we used $\|\tilde{K}\|\leq f(\lambda)$. Therefore
$\mathrm{tr}(W_{\epsilon})\leq\frac{1}{1-2\mathrm{tr}(S^{\star})f(\lambda)\|B\|}\mathrm{tr}(W_{\epsilon}^{0}).$ (17)
Combining equations (16)-(17) yields
$\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})\leq\frac{\mathrm{tr}((W_{\epsilon}^{0})^{-1}V^{-1})}{1{-}\frac{2\kappa(V)f(\lambda)\left\lVert V\right\rVert\left\lVert B\right\rVert\|U^{\star}\|\mathrm{tr}(W_{\epsilon}^{0})\mathrm{tr}((W^{0}_{\epsilon})^{-1}V^{-1})}{1-2\mathrm{tr}(S^{\star})f(\lambda)\|B\|}},$
which is equivalent to the final result.
Note that $f(\lambda)\sim O(\lambda)$ for large $\lambda$ and $f(\lambda)\sim O(\sqrt{\lambda})$ for small $\lambda$. Looking at (12), this implies that small performance losses $\lambda$ yield steeper trade-offs in reducing adversarial observability for the cost (12). On the other hand, the lower bound in Problem 2 can become unbounded as $\lambda$ increases, since a larger $\lambda$ allows the Gramian inverse to become singular. As a result, higher values of $\lambda$ can produce sharper improvements in the objective of Problem 2. An illustration of the function $f(\lambda)$ is provided in Figure 1 for the first example of the simulations in Section 5.
The lower bounds also depend on a number of other parameters. Most notable is $\left\lVert B\right\rVert$, which dictates that larger actuation authority for the operator allows for sharper decrease in adversarial observability. They also depend on $V$, which captures the interplay between observability minimization and initial condition covariance.
Next, note that although the lower bound of Problem 2 is local, the objective value of Problem 2 remains globally bounded due to the regularization parameter $\epsilon$. Consequently, we can extend the lower bound to hold globally as follows.
Proposition 1.
For all $\lambda\geq 0$:
$J_{2}(\lambda)\geq-\frac{2\mathrm{tr}(V^{-1})(\left\lVert A+BK^{\star}\right\rVert+\left\lVert B\right\rVert f(\lambda))}{\epsilon}$
Proof.
From the second constraint of Problem 2 we obtain
$W_{\epsilon}$ $=\int_{0}^{\infty}e^{(A+BK)^{\mathrm{T}}t}(C^{\mathrm{T}}C+\epsilon I)e^{(A+BK)t}\textrm{d}t$
$\geq\epsilon\int_{0}^{\infty}e^{(A+BK)^{\mathrm{T}}t}e^{(A+BK)t}\textrm{d}t$
$\geq\epsilon\int_{0}^{\infty}e^{-2\left\lVert A+BK\right\rVert t}\textrm{d}t=\frac{\epsilon}{2\left\lVert A+BK\right\rVert},$
hence $\underline{\lambda}(W_{\epsilon})\geq\frac{\epsilon}{2\left\lVert A+BK\right\rVert}$. Therefore:
$\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})\leq\mathrm{tr}(V^{-1})\bar{\lambda}(W_{\epsilon}^{-1})\\ =\frac{\mathrm{tr}(V^{-1})}{\underline{\lambda}(W_{\epsilon})}\leq\frac{2\mathrm{tr}(V^{-1})\left\lVert A+BK\right\rVert}{\epsilon}.$ (18)
Moreover, from Lemma 1:
$\left\lVert A{+}BK\right\rVert\leq\|A{+}BK^{\star}{+}B\tilde{K}\|\leq\left\lVert A{+}BK^{\star}\right\rVert+\left\lVert B\right\rVert f(\lambda).$ (19)
Combining (18)-(19) yields the required result.
Combining Theorem 1 with Proposition 1 yields lower bounds for Problems 1 and 2 across all values of the trade-off parameter $\lambda$ that quantify the fundamental limits of adversarial observability reduction under a given level of performance degradation.
Figure 1: Illustration of the function $f(\lambda)$.
3.2 Robustness to Adversary’s Sensing Uncertainty
Assumption 2 requires knowledge of the adversary’s sensing matrix $C$. This is often realistic in practice, for instance, when the adversary employs radar with known location and measurement geometry to track an aircraft. In some settings, however, the system operator may only know a superset of the adversary’s possible sensing configurations, rather than the exact matrix $C$.
Solving Problems 1 and 2 using such a superset generally yields conservative (and thus suboptimal) solutions. Nevertheless, any controller obtained in this manner provides a certified upper bound on the adversary’s observability metric, as we establish next.
Proposition 2.
Assume that the adversary’s sensing matrix is $\hat{C}\in\mathbb{R}^{\hat{m}\times n}$, where $\hat{m}\leq m$, and where the rows of $\hat{C}$ are a subset of the rows of $C$. Then, $J_{oj}(K;~\hat{C})\leq J_{oj}(K;~C)$, $j=1,2$.
Proof.
Denote as $W_{C}$ and $W_{\hat{C}}$ the observability Gramians corresponding to $C$ and $\hat{C}$. Then, by definition, we have $(A+BK)^{\mathrm{T}}W_{C}+W_{C}(A+BK)+C^{\mathrm{T}}C=0$ and $(A+BK)^{\mathrm{T}}W_{\hat{C}}+W_{\hat{C}}(A+BK)+\hat{C}^{\mathrm{T}}\hat{C}=0$. Since $\hat{C}^{\textrm{T}}\hat{C}=C^{\textrm{T}}S^{\textrm{T}}SC$ for a row selection matrix $S\in\mathbb{R}^{\hat{m}\times m}$, we have $C^{\textrm{T}}C-\hat{C}^{\textrm{T}}\hat{C}=C^{\textrm{T}}(I-S^{\textrm{T}}S)C\succeq 0$. Hence, $W_{\hat{C}}\preceq W_{C}$, and thus $J_{o1}(K;~\hat{C})=\textrm{tr}(W_{\hat{C}}V)\leq\textrm{tr}(W_{C}V)=J_{o1}(K;~C)$. The fact that $J_{o2}(K;~\hat{C})\leq J_{o2}(K;~C)$ follows similarly since negated inversion preserves matrix ordering.
Another possible sensing uncertainty concerns perturbations in the numerical values of $C$, rather than uncertainty in its rows (e.g., $\hat{C}=C+\delta C$ for some small $\delta C$). In such cases, Problem 1 is inherently more robust since its objective depends quadratically on $C$ (see next section), whereas Problem 2 can be more sensitive to perturbations in $C$, particularly when the Gramian approaches singularity due to the inverse-based objective. One could also address such uncertainty in the worst-case sense through a robust SDP formulation, though this remains beyond the scope of this work.
4 Algorithmic Solutions to Problems 1 and 2
We subsequently provide algorithmic solutions to Problems 1 and 2. Specifically, we provide an exact algorithmic solution for Problem 1 by reformulating it into an equivalent SDP, and an approximate algorithmic solution for Problem 2 using sequential SDP.
4.1 Solution to Problem 1 using an SDP
We begin by showing that Problem 1 can be cast as an SDP.
Theorem 2.
The solution to Problem 1 is given by $K=XS^{-1}$, where $X$ and $S$ solve the SDP
$\begin{split}\min_{X,S,Z}\quad&\mathrm{tr}(CSC^{\mathrm{T}})\\ \mathrm{s.t.}\quad&AS+BX+SA^{\mathrm{T}}+X^{\mathrm{T}}B^{\mathrm{T}}+V=0,\\ &\mathrm{tr}(SQ)+\mathrm{tr}(Z)\leq\mathrm{tr}(P^{\star}V)+\lambda,\\ &\begin{bmatrix}Z&R^{1/2}X\\ X^{\mathrm{T}}R^{1/2}&S\end{bmatrix}\succeq 0,~Z\succeq 0,~S\succeq 0.\end{split}$ (20)
Proof.
By the constraints of Problem 1, we have $W=\int_{0}^{\infty}e^{(A+BK)^{\textrm{T}}\tau}C^{\mathrm{T}}Ce^{(A+BK)\tau}\textrm{d}\tau$ and $P=\int_{0}^{\infty}e^{(A+BK)^{\mathrm{T}}\tau}(Q+K^{\mathrm{T}}RK)e^{(A+BK)\tau}\textrm{d}\tau$. Therefore, using the cyclic property of the trace operator, we obtain $\mathrm{tr}(WV)=\mathrm{tr}(CSC^{\mathrm{T}})$ and $\mathrm{tr}(PV)=\mathrm{tr}(S(Q+K^{\mathrm{T}}QK))$, where $S\succ 0$ is the unique solution of the Lyapunov equation $(A+BK)S+S(A+BK)^{\mathrm{T}}+V=0.$ Using these equivalences, we can reduce the dimension of Problem 1 and conclude it is equivalent to the program
$\begin{split}\min_{K,S}\quad&\mathrm{tr}(CSC^{\mathrm{T}})\\ \textrm{s.t.}\quad&(A+BK)S+S(A+BK)^{\mathrm{T}}+V=0,\\ &\mathrm{tr}(S(Q+K^{\mathrm{T}}RK))\leq\mathrm{tr}(P^{\star}V)+\lambda,~S\succeq 0.\end{split}$ (21)
Performing the change of variables $X=KS\Longrightarrow K=XS^{-1}$, it follows that (21) has the same optimal solution as
$\begin{split}\min_{X,S}\quad&\mathrm{tr}(CSC^{\mathrm{T}})\\ \textrm{s.t.}\quad&AS+BX+SA^{\mathrm{T}}+X^{\mathrm{T}}B^{\mathrm{T}}+V=0,~S\succeq 0,\\ &\mathrm{tr}(SQ)+\mathrm{tr}(R^{1/2}XS^{-1}X^{\mathrm{T}}R^{1/2})\leq\mathrm{tr}(P^{\star}V)+\lambda.\end{split}$ (22)
Introducing the variable $Z\succeq R^{1/2}XS^{-1}X^{\mathrm{T}}R^{1/2}$ preserves the optimal solution. Therefore, (22) is equivalent to
$\begin{split}\min_{X,S,Z}\quad&\mathrm{tr}(CSC^{\mathrm{T}})\\ \textrm{s.t.}\quad&AS+BX+SA^{\mathrm{T}}+X^{\mathrm{T}}B^{\mathrm{T}}+V=0,\\ &\mathrm{tr}(SQ)+\mathrm{tr}(Z)\leq\mathrm{tr}(P^{\star}V)+\lambda,\\ &Z-R^{1/2}XS^{-1}X^{\mathrm{T}}R^{1/2}\succeq 0,Z\succeq 0,S\succeq 0.\end{split}$ (23)
Finally, using the Schur complement, we obtain the SDP (20).
Theorem 2 enables us to reformulate the original Problem 1 as an SDP where both the objective and constraints are convex. This convex formulation preserves the structural properties of the original problem while allowing for efficient computation using off-the-shelf SDP solvers. Algorithm 1 outlines the resulting solution procedure.
4.2 Approximate Solution to Problem 2 using Sequential SDP
Algorithm 1 SDP for Problem 1
Output: Exact solution $(K,P,W)$ to Problem 1.
1:procedure
2:  Solve for $(X,S,Z)$ from (20).
3:  Compute $K=XS^{-1}$.
4:  Compute P,WP,~W from the equality constraints of Problem 1.
5:end procedure
Since the negated trace of a matrix inverse is concave in the cone of positive definite matrices, Problem 2 is not convex and thus cannot be cast as an SDP. However, we can formulate an algorithm to approximately solve it using the convex-concave procedure.
Toward this end, one challenge entailed in Problem 2 is that its cost function is nonlinear and, most importantly, it decreases rapidly as the matrix $W_{\epsilon}$ approaches singularity. As a result, linearizing this cost can lead to large approximation errors that slow down the convergence of a convex-concave procedure. Nevertheless, by transforming the Lyapunov equation for the observability Gramian into a special form of an algebraic Riccati equation, we can effectively deal with this issue. Specifically, we show Problem 2 can be cast as an equivalent difference of convex functions (DC) program with a linear objective function that involves no inverses of matrix variables.
Lemma 2.
Under the transformation $Y_{\epsilon}=W_{\epsilon}^{-1}$, Problem 2 admits the same optimal solution as the DC program with linear cost:
$\min_{K,P,Y_{\epsilon}}$ $\quad-\mathrm{tr}(Y_{\epsilon}V^{-1})$ (24)
$\mathrm{s.t.}\quad$ $\begin{bmatrix}\frac{1}{2}(A+BK+P)^{\mathrm{T}}(A+BK+P)+Q&K^{\mathrm{T}}\\ \star&-R^{-1}\end{bmatrix}$
$-\begin{bmatrix}\frac{1}{2}(A+BK-P)^{\mathrm{T}}(A+BK-P)&0\\ \star&0\end{bmatrix}\preceq 0,$
$\begin{bmatrix}\frac{1}{2}(A+BK+Y_{\epsilon})(A+BK+Y_{\epsilon})^{\mathrm{T}}&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C{+}\epsilon I}\\ \star&I\end{bmatrix}$
$-\begin{bmatrix}\frac{1}{2}(A+BK-Y_{\epsilon})(A+BK-Y_{\epsilon})^{\mathrm{T}}&0\\ \star&0\end{bmatrix}\preceq 0,$
$\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~Y_{\epsilon}\succeq 0.$
Proof.
We first rewrite the cost function of Problem 2 into a linear one. To that end, note that since $P\succeq 0$ and $Q+K^{\mathrm{T}}RK\succ 0$, the first constraint of Problem 2 is a Lyapunov equation for $A+BK$, hence $A+BK$ is Hurwitz [3]. Therefore, the second constraint in Problem 2 implies $W_{\epsilon}\succ 0$ strictly since $C^{\textrm{T}}C+\epsilon I\succ 0$, hence $W_{\epsilon}^{-1}$ exists. In light of this, multiplying the second constraint of Problem 2 from the left and from the right with $W_{\epsilon}^{-1}$, we obtain the equation
$Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0,$ (25)
for which $Y_{\epsilon}=W_{\epsilon}^{-1}$ is a solution. However, (25) also admits other positive semidefinite solutions. We thus need to show that using (25) in place of the second constraint in Problem 2 preserves optimality. To this end, note that we can rewrite (25) as
$\hskip-8.53581pt-Y_{\epsilon}(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))^{\mathrm{T}}-(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))Y_{\epsilon}\\ +Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0.$
Since $C^{\mathrm{T}}C+\epsilon I\succ 0$, for $Y_{\epsilon}=W_{\epsilon}^{-1}\succ 0$ we have $Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}\succ 0$. By the Lyapunov Theorem, $(A+BK+Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I))^{\mathrm{T}}$ is strictly stable for $Y_{\epsilon}=W_{\epsilon}^{-1}$, hence $Y_{\epsilon}=W_{\epsilon}^{-1}$ is the stabilizing solution to (25), and hence also the maximal one by Theorem 12.18.4 in [3]. Therefore, for any other solution $\hat{Y}_{\epsilon}\neq W_{\epsilon}^{-1}$ to (25) we have $\hat{Y}_{\epsilon}\prec W_{\epsilon}^{-1}$ and thus $\mathrm{tr}(\hat{Y}_{\epsilon}V^{-1})>-\mathrm{tr}(W_{\epsilon}^{-1}V^{-1})$. It thus follows that if we substitute the second constraint in Problem 2 with (25) and perform the variable change $Y_{\epsilon}=W_{\epsilon}^{-1}$, the optimal solution will remain the same. Hence, Problem 2 is equivalent to the following optimization problem with linear cost:
$\min_{K,P,Y_{\epsilon}}$ $\quad-\mathrm{tr}(Y_{\epsilon}V^{-1})$ (26)
s.t. $(A+BK)^{\mathrm{T}}P+P(A+BK)+Q+K^{\mathrm{T}}RK=0,$
$-Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}=0,$
$\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~Y_{\epsilon}\succeq 0.$
Subsequently, it is straightforward that relaxing the first equality constraint in (26) into a less-than-equal-to inequality preserves the optimal solution and the stability of $A+BK$. In addition, since $\mathrm{tr}(Y_{\epsilon}V^{-1})$ appears negatively in the cost function, we can also relax the second constraint in (26) into a greater-than-equal-to inequality. Employing these relaxations and using the Schur complement, it follows that (26) is equivalent to
$\min_{K,P,Y_{\epsilon}}$ $\quad-\mathrm{tr}(Y_{\epsilon}V^{-1})$
s.t. $\begin{bmatrix}(A+BK)^{\mathrm{T}}P+P(A+BK)+Q&K^{\mathrm{T}}\\ \star&-R^{-1}\end{bmatrix}\preceq 0,$
$\begin{bmatrix}Y_{\epsilon}(A+BK)^{\mathrm{T}}+(A+BK)Y_{\epsilon}&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C{+}\epsilon I}\\ \star&-I\end{bmatrix}\preceq 0,$
$\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~Y_{\epsilon}\succeq 0.$
The final result follows by using $Y_{\epsilon}(A+BK)^{\mathrm{T}}+(A+BK)Y_{\epsilon}=\frac{1}{2}(A+BK+Y_{\epsilon})(A+BK+Y_{\epsilon})^{\mathrm{T}}-\frac{1}{2}(A+BK-Y_{\epsilon})(A+BK-Y_{\epsilon})^{\mathrm{T}}$ and $(A+BK)^{\mathrm{T}}P+P(A+BK)=\frac{1}{2}(A+BK+P)^{\mathrm{T}}(A+BK+P)-\frac{1}{2}(A+BK-P)^{\mathrm{T}}(A+BK-P)$.
Let us now define the functions [^1] We use the subscript $j$ with a slight abuse of notation.:
$L_{j}(X,Z)$ $=\frac{1}{2}(A+BX_{j}-Z_{j})^{\mathrm{T}}(A+BX_{j}-Z_{j})$
$+\frac{1}{2}(A+BX_{j}-Z_{j})^{\mathrm{T}}(B(X-X_{j})-(Z-Z_{j}))$
$+\frac{1}{2}(B(X-X_{j})-(Z-Z_{j}))^{\mathrm{T}}(A+BX_{j}-X_{j}),$
$L_{j}^{\prime}(X,Z)$ $=\frac{1}{2}(A+BX_{j}-Z_{j})(A+BX_{j}-Z_{j})^{\mathrm{T}}$
$+\frac{1}{2}(A+BX_{j}-Z_{j})(B(X-X_{j})-(Z-Z_{j}))^{\mathrm{T}}$
$+\frac{1}{2}(B(X-X_{j})-(Z-Z_{j}))(A+BX_{j}-X_{j})^{\mathrm{T}}.$
To use the convex-concave procedure and obtain an approximate solution to Problem 2, we linearize the concave parts of the constraints of (24) about a point $(K_{j},P_{j},W_{j})$ and obtain
$\min_{K,P,Y_{\epsilon}}$ $~~-\mathrm{tr}(Y_{\epsilon}V^{-1})$
$\mathrm{s.t.}\quad$ $\hskip-5.69054pt\begin{bmatrix}\frac{1}{2}(A{+}BK{+}P)^{\mathrm{T}}(A{+}BK{+}P){+}Q{-}L_{j}(K,P)&K^{\mathrm{T}}\\ \star&-R^{-1}\end{bmatrix}{\preceq}0,$
$\hskip-25.60747pt\begin{bmatrix}\frac{1}{2}(A{+}BK{+}Y_{\epsilon})(A{+}BK{+}Y_{\epsilon})^{\mathrm{T}}{-}L_{j}^{\prime}(K,Y_{\epsilon})&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C{+}\epsilon I}\\ \star&I\end{bmatrix}{\preceq}0,$
$\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~Y_{\epsilon}\succeq 0.$
Using the Schur complement, this is equivalent to
$\min_{K,P,Y_{\epsilon}}$ $\quad-\mathrm{tr}(Y_{\epsilon}V^{-1})$ (27)
$\mathrm{s.t.}\quad$ $\begin{bmatrix}Q-L_{j}(K,P)&K^{\mathrm{T}}&\frac{1}{\sqrt{2}}(A+BK+P)^{\mathrm{T}}\\ \star&-R^{-1}&0\\ \star&\star&-I\end{bmatrix}\preceq 0,$
$\hskip-17.07164pt\begin{bmatrix}-L_{j}^{\prime}(K,Y_{\epsilon})&Y_{\epsilon}\sqrt{C^{\mathrm{T}}C+\epsilon I}&\frac{1}{\sqrt{2}}(A+BK+Y_{\epsilon})\\ \star&-I&0\\ \star&\star&-I\end{bmatrix}\preceq 0,$
$\mathrm{tr}((P-P^{\star})V)\leq\lambda,~P\succeq 0,~Y_{\epsilon}\succeq 0,$
the iterative solution of which, as shown in in Algorithm 2, should eventually lead us to a stationary point of (24).
Remark 5.
Algorithm 2 requires initialization with sets of matrices $K_{0},P_{0},Y_{\epsilon,0}$ that are feasible for Problem 2. One such initialization is obtained by selecting $K_{0}$ as the linear-quadratic gain corresponding to $\lambda=0$. The matrices $P_{0}$ and $Y_{\epsilon,0}$ are then obtained by solving the first and second Lyapunov equations of Problem 2 under $K=K_{0}$, respectively, with $Y_{\epsilon,0}$ computed as the inverse of the resulting $W_{\epsilon}$. With such an initialization, (27) will also be feasible.
Algorithm 2 Sequential SDP for Problem 2
Input: Feasible point $K_{0},P_{0},Y_{\epsilon,0}$ to Problem 2, tolerance $\delta>0$.
   Output: Estimated solution $(\hat{K},\hat{P},\hat{Y}_{\epsilon})$ to Problem 2.
1:procedure
2:  $j\leftarrow 0$.
3:  while $j=0$ or $|\mathrm{tr}(Y_{\epsilon,j}-Y_{\epsilon,j-1})|\geq\delta$ do
4:   Solve for $K,P,Y_{\epsilon}$ from (27).
5:   $(K_{j+1},P_{j+1},Y_{\epsilon,j+1})\leftarrow(K,P,Y_{\epsilon})$.
6:   $j\leftarrow j+1$.
7:  end while
8:  $(\hat{K},\hat{P},\hat{Y}_{\epsilon})\leftarrow(K,P,Y_{\epsilon})$.
9:end procedure
We summarize the convergence properties of Algorithm 2 in the following theorem. This convergence relies fundamentally on the regularization of Problem 2 with the parameter $\epsilon>0$.
Theorem 3.
Algorithm 2 converges to a stationary point of (24).
Proof.
Note that Problem 2 is equivalent to (24) according to Lemma 2. In addition, (24) is a DC program wherein all functions involved are continuously differentiable with respect to $K,P,Y_{\epsilon}$. Moreover, by Lemma 1 and the constraint $\textrm{tr}((P-P^{\star})V)\leq\lambda$, boundedness of the feasible set for $K$ and $P$ follows. Finally, note that the second constraint in (24) implies
$Y_{\epsilon}(A+BK)^{\mathrm{T}}-(A+BK)Y_{\epsilon}-Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}\succeq 0.$ (28)
Since the feasible set for $K$ is bounded, (A+BK)-(A+BK) is bounded. Therefore, if $Y_{\epsilon}\succeq 0$ and $\left\lVert Y_{\epsilon}\right\rVert\rightarrow\infty$, then (28) cannot hold because the quadratic term $Y_{\epsilon}(C^{\mathrm{T}}C+\epsilon I)Y_{\epsilon}$ will force the left-hand side of (28) to become negative definite. This is particularly true because $C^{\textrm{T}}C+\epsilon I\succ 0$, following $\epsilon>0$. Hence the feasible set for $Y_{\epsilon}$ is bounded above, hence for any $X\succeq 0$ the set of feasible $Y_{\epsilon}$ satisfying $\mathrm{tr}(Y_{\epsilon}V^{-1})\leq-\mathrm{tr}(XV^{-1})$ is bounded. Therefore, by [18, 17], Algorithm 2 converges to a stationary point of (24).
Remark 6.
By Lemma 2, (24) has the same optimal solution as Problem 2. Thus, Theorem 3 ensures that if Algorithm 2 converges to a minimum, this will also be the minimum of Problem 2.
5 Numerical Examples
5.1 Intuitive Example
To provide an easy understanding of our algorithm, we first run simulations on the following system:
$\dot{x}(t)$ $=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}x(t)+\begin{bmatrix}1\\ 1\end{bmatrix}u(t),\quad y(t)=\begin{bmatrix}1&0\end{bmatrix}x(t).$
This system has the interesting property that it is observable from $C=\begin{bmatrix}1&0\end{bmatrix}$, but at the same time, by inspection, we can infer that it becomes unobservable with the state feedback $u=-x_{2}$. This feedback clearly decouples the first state from the second one, making the latter unobservable. Still, here we are interested in a feedback controller that not only makes the system less observable from $C$, but also has a hard upper bound on its distance from optimality.
Figure 2: Minimum values of Problems 1 and 2 as computed by Algorithm 1 and 2, as well as the corresponding lower bounds calculated in Section 3, for various values of the trade-off parameter $\lambda$.
To get a controller that balances loss observability from $C$ with optimal performance, we run Algorithm 2 with parameters $Q=0.2I_{2}$, $R=1$, $\delta=10^{-3}$, $\epsilon=10^{-4}$, $V=I_{2}$, and initialize $K_{0}=K^{\star}$ based on the positive definite solution of the ARE (6). To solve the underlying SDPs, we use CVX, a package for specifying and solving convex programs [9, 11]. With $\lambda=0.01$ and $\lambda=0.1$, the closed-loop matrices $A+B\hat{K}$ that we obtain based on the output of Algorithm 2 are
$\lambda=0.01:~~A+B\hat{K}$ $=\begin{bmatrix}-0.4873&0.1905\\ -0.4873&-0.8095\end{bmatrix}$,
$\lambda=0.1:~~A+B\hat{K}$ $=\begin{bmatrix}-0.8572&-0.0018\\ -0.8572&-1.0018\end{bmatrix}$,
whereas the nominal optimal closed-loop matrix $A+BK^{\star}$ is:
$A+BK^{\star}=\begin{bmatrix}-0.4472&0.3095\\ -0.4472&-0.6905\end{bmatrix}$.
The pattern we notice is the one we would ideally expect: as we gain more flexibility to deviate from strict optimality, the resulting control gain $\hat{K}$ increasingly decouples the first state of the system from the second one. This is achieved primarily by forcing the upper-right entry of the matrix $A+B\hat{K}$ to approach zero. In doing so, we obtain a control policy that remains close to optimal in terms of performance, while simultaneously making it significantly harder to observe the system through the output matrix $C$.
Figure 2 shows the trade-off between performance and adversarial observability. Specifically, it shows the minimum values of Problems 1 and 2 as computed by Algorithm 1 and 2, as well as the corresponding lower bounds we calculated in Section 3, for various values of the trade-off parameter $\lambda$. We verify that the theoretically derived lower bounds are correct. While they are relatively tight for Problem 1, they become looser for Problem 2 because the inverse of the Gramian tends to be close to singular as $\lambda$ increases, and because the obtained solution is only a local minimum. Finally, we also validate the intuition behind Theorem 3: when minimizing the trace of the observability Gramian, smaller performance trade-offs yield a larger payoff in adversarial observability. On the other hand, when maximizing the trace of the Gramian’s inverse, the payoff is steeper when the performance trade-off becomes larger.
5.2 ADMIRE Aircraft
We perform further simulations on the ADMIRE aircraft [15], which has $n=5$ states and $m=7$ control inputs. We assume that an adversary can measure the first, third, and fifth state, while we refer to [15] for the expressions of the matrices $A,B$.
Figure 3: Evolution of the eigenvalues $\lambda_{i}(W)$, $i=1,\ldots,5$, of the observability Gramian $W$ and of the trace of its inverse, $\mathrm{tr}(W^{-1})$, for $(A+BK_{j},C)$ during Algorithm 2.
Table 1: Eigenvalues under the nominal optimal gain $K^{\star}$ and the gain obtained by Algorithms 1-2, and the corresponding costs.
$\textrm{tr}(W)$ $\textrm{tr}(W^{-1})$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$
Nom. $13.82$ $558$ $9.79$ $3.63$ $0.341$ $0.0709$ $0.00180$
Alg. 1 $11.01$ $598$ $9.34$ $1.28$ $0.331$ $0.0520$ $0.00180$
Alg. 2 $13.80$ $31504$ $9.78$ $3.60$ $0.351$ $0.0707$ $0.00003$
First, we perform Algorithm 2, to obtain a control gain $\hat{K}$ that balances the loss of observability from $C$ with the performance of the closed-loop system $A+B\hat{K}$. We choose the parameters of the algorithm as $Q=I_{5}$, $R=10I_{7}$, $\delta=10$ (scaled according to this setup’s cost function), $\epsilon=10^{-5}$, $V=I_{5}$, $\lambda=1$, and initialize $K_{0}=K^{\star}$ based on the solution of the ARE (6).
Figure 3 shows the evolution of the eigenvalues – and the trace of the inverse – of the observability Gramian of $(A+BK_{j},C)$ throughout the execution of Algorithm 2. We observe that the eigenvalue that is affected the most is the minimum one, which is driven very close to zero (see also Table 1). This is both an expected and a desirable property. On the one hand, maximizing the trace of the inverse of the observability Gramian is equivalent to making the Gramian singular, which is indeed most easily achieved by making the minimum eigenvalue equal to zero. On the other hand, this outcome means that the resulting closed-loop pair $(A+B\hat{K},C)$ is very close to being unobservable, and hence an adversary who knows $\hat{K}$ and who observes $y(t)$ will not be able to easily reconstruct the full state $x(t)$ of the system. The figure also illustrates a substantial increase in the trace of the observability Gramian, all while maintaining acceptable control performance — as measured by the linear-quadratic cost — with a relatively small compromise, limited to a trade-off of $\lambda=1$.
Second, we perform Algorithm 1, to obtain a control gain $K$ that minimizes observability of $(A+BK,C)$ on average while trading off with performance. We use the same parameter values as in the previous example. Table 1 shows the eigenvalues and the trace of the observability Gramian both under the optimal gain $K^{\star}$ and under the gain obtained from Algorithm 1. The pattern here differs from that in the case of Algorithm 2; instead of trying to minimize the minimum eigenvalue as much as possible, the algorithm focuses more on minimizing the larger eigenvalues of the Gramian. This is also an expected behavior, as there is significantly more to be gained in terms of the cost function by minimizing the largest eigenvalues, instead of the minimum eigenvalue that is many orders of magnitude smaller. Another noteworthy detail is that, out of the five eigenvalues of the Gramian, only the second and the fourth one underwent significant relative reduction. This is owed to the fact that the adversary can observe three out of the five states, so only two observability eigenvalues can decrease significantly at a time.
(a) Trajectories under the optimal feedback gain.
(b) Trajectories under the feedback gain from Algorithm 1.
(c) Trajectories under the feedback gain from Algorithm 2.
Figure 4: Trajectories of the adversary’s observation errors, and the closed-loop performance cost.
Finally, to showcase the adversarial observability and performance trade-offs provided by Algorithms 1 and 2, we simulate the closed-loop behavior of the system along with a Luenberger observer – used by an adversary to reconstruct the system state – that places observation poles to 15;5;2;2;1-15;-5;-2;-2;-1. For this state estimation scenario, we assume the sensing components of the adversary are corrupted by deterministic noise composed of $5$ sinusoids, with frequencies between 0 to $\frac{1}{2\pi}$ and magnitudes equal to $0.01$. Figure 4 shows the incurred closed-loop performance cost along with the estimation errors of the adversary under a) the nominal optimal controller; b) the controller obtained by Algorithm 1; and c) the controller obtained by Algorithm 2. We observe that the controller for Algorithm 1 magnifies adversarial estimation errors across all states with almost negligible loss of performance, while Algorithm 1 increases the estimation error of the fourth state to completely unreliable levels with a little more performance trade-off. This pattern also aligns with the interpretation of the metrics (4): on the one hand, the trace of the Gramian captures estimation capabilities across all directions in the state space; on the other hand, the trace of the Gramian inverse is skewed toward the least observable state which, in this example, corresponds to the fourth state of the aircraft.
6 Conclusion
We study the problem of feedback control that minimizes the observability of a set of sensors deployed by an adversary, while conforming to strict closed-loop performance constraints. We quantify observability using metrics related to the trace of the observability Gramian and its inverse. For both of these, we establish theoretical lower bounds on their optimal values as a function of the performance trade-off parameter. Finally, we optimize the trace of the observability Gramian through an SDP, while, for the trace of the Gramian inverse, we perform this optimization approximately using sequential SDP.
Future work includes the consideration of strict output feedback controllers that rely only on partial state information, but which still achieve a balance between having sufficient performance and minimizing adversarial observability.